# How Far Are We From AGI?

### v1

@@ 7.6 Embodied AI: AI for Robotics
Novel AI Application in Recent Robotics Research Wake et al. (2023) proposes a novel pipeline that enhances GPT-4V(vision), a general-purpose Vision Language Model, by integrating observations of human actions to facilitate robotic manipulation. Yell At Your Robot (YAY Robot) system (Shi et al., 2024) allows robots to adapt to verbal corrections in real-time and improve upon their high-level policy decisions iteratively. This system leverages Language-Conditioned Behavior Cloning (LCBC) to learn a wide range of skills specified through language, enabling users to interact with robots using free-form commands. Zhang et al. (2023f) introduces the NOIR system, an innovative brain-robot interface (BRI) that employs non-invasive electroencephalography (EEG) to enable humans to command robots to perform a diverse range of everyday activities.
@@ 8 Conclusion
In addition to establishing a shared groundwork for AI researchers through a comprehensive examination of the latest research advancements, we also articulate our vision of the fundamental nature of AGI and advocate for a responsible approach to its development. Our goal here is to offer concrete directions for further exploration and to spark robust, thought-provoking discussions that will advance the community toward the realization of “true” AGI. Given the continually evolving definition and objectives of AGI research, we intend to regularly update this manuscript to incorporate fresh insights and breakthroughs from the research community. Note that the visions we present are inherently limited and incomplete. Our objective is to stimulate brainstorming within the AI community, and we eagerly await the emergence of superior visions from within the community itself.
@@
The pursuit of AGI is an extraordinarily challenging endeavor, fraught with immense complexities and obstacles. However, with the recent advancements in various aspects of AI research, the vision of achieving AGI has never been more vivid and tangible. Prominent figures in the AI community, such as Jensen Huang  [^21] https://www.thestreet.com/technology/nvidia-ceo-jensen-huang-artificial-general-intelligence, Sam Altman  [^22] https://time.com/6344160/a-year-in-time-ceo-interview-sam-altman/, and Elon Musk  [^23] https://www.reuters.com/technology/teslas-musk-predicts-ai-will-be-smarter-than-smartest-human-next-year-2024-04-08/, have expressed their confidence in the eventual realization of AGI, further fueling the enthusiasm and determination of researchers and engineers worldwide.

### v2

@@ 7.6 Embodied AI: AI for Robotics
Novel AI Application in Recent Robotics Research Wake et al. (2023) proposes a novel pipeline that enhances GPT-4V(vision), a general-purpose Vision Language Model, by integrating observations of human actions to facilitate robotic manipulation. Yell At Your Robot (YAY Robot) system (Shi et al., 2024a) allows robots to adapt to verbal corrections in real time and improve upon their high-level policy decisions iteratively. This system leverages Language-Conditioned Behavior Cloning (LCBC) to learn a wide range of skills specified through language, enabling users to interact with robots using free-form commands. Zhang et al. (2023f) introduces the NOIR system, an innovative brain-robot interface (BRI) that employs non-invasive electroencephalography (EEG) to enable humans to command robots to perform a diverse range of everyday activities.
@@ 8 Conclusion
In addition to establishing a shared groundwork for AI researchers through a comprehensive examination of the latest research advancements, we also articulate our vision of the fundamental nature of AGI and advocate for a responsible approach to its development. Our goal here is to offer concrete directions for further exploration and to spark robust, thought-provoking discussions that will advance the community toward the realization of “true” AGI. Given the continually evolving definition and objectives of AGI research, we intend to regularly update this manuscript to incorporate fresh insights and breakthroughs from the research community. Note that the visions we present are inherently limited and incomplete. Our objective is to stimulate brainstorming within the AI community, and we eagerly await the emergence of superior visions from within the community itself. Here are the main contributions of our work:
@@
Acknowledgments
We would like to sincerely thank the speakers of the “How Far Are We from AGI” workshop, including Oriol Vinyals, Yejin Choi, Andrew Gordon Wilson, Song Han, and Yoshua Bengio, for bringing many insights to the discussion of this survey paper. We also thank Haofei Yu, Jinwei Yao, Zhong Li, Pengju Yan, Weihao Tan, and Tianmin Shu for providing helpful feedback on our manuscript and GitHub repository.